# Volumetric medical image segmentation through dual self-distillation in U-shaped networks

### v2

@@ Abstract
U-shaped networks and its variants have demonstrated exceptional results for medical image segmentation. In this paper, we propose a novel dual self-distillation (DSD) framework in U-shaped networks for volumetric medical image segmentation. DSD distills knowledge from the ground-truth segmentation labels to the decoder layers. Additionally, DSD also distills knowledge from the deepest decoder and encoder layer to the shallower decoder and encoder layers respectively of a single U-shaped network. DSD is a generalized training strategy that could be attached to the backbone architecture of any U-shaped network to further improve its segmentation performance. We attached DSD on several state-of-the-art U-shaped backbones, and extensive experiments on various public 3D medical image segmentation datasets (cardiac substructure, brain tumor and Hippocampus) demonstrated significant improvement over the same backbones without DSD. On average, after attaching DSD to the U-shaped backbones, we observed an increase of 2.82%, 4.53% and 1.3% in Dice similarity score, a decrease of 7.15 mm, 6.48 mm and 0.76 mm in the Hausdorff distance, for cardiac substructure, brain tumor and Hippocampus segmentation, respectively. These improvements were achieved with negligible increase in the number of trainable parameters and training time. Our proposed DSD framework also led to significant qualitative improvements for cardiac substructure, brain tumor and Hippocampus segmentation over the U-shaped backbones. The source code is publicly available at https://github.com/soumbane/DualSelfDistillation.
@@ 1 Introduction
The need of transferring knowledge from a large teacher network to a smaller student network was later eliminated by the process of self-distillation [30]. Self-distillation uses the deepest layer of a single model to act as the teacher network in order to distill knowledge to the shallower layers of the same model, acting as the student network. Since self-distillation uses a single model, it is much more efficient when compared to ensemble methods. Additionally, self-distillation is much faster to train as it is not necessary to train a large network as is required by knowledge distillation. Self-distillation has been applied for numerous computer vision tasks such as image classification [30, 35] and object detection [34]. A rigorous theoretical explanation of ensemble, knowledge distillation and self-distillation for image classification was recently provided in [47]. Recently, the self-distillation process was applied to a transformer based TransUNet [49], along with multi-scale fusion blocks to perform 3D medical image segmentation [36]. However, this method was limited to the specific TransUNet architecture.
@@
In this paper, we address the question of whether we can train a single U-shaped network and leverage the learned representations of the deeper encoder and decoder network layers to improve the prediction performance of medical image segmentation. Here, we propose a novel volumetric dual self-distillation (DSD) framework that could independently be attached to any U-shaped image segmentation backbone. In DSD, the deepest encoder and decoder of the U-shaped backbone act as the teacher networks for the shallower encoders and decoders which act as student networks. We found that the deepest encoder which is at the bottom of the contracting encoder path of the U-shaped network contains more contextual information compared to the shallower encoder layers. Similarly, the deepest decoder which is at the top of the expanding decoder path of the U-shaped network contains more semantic information than the shallower decoder layers. Thus, in our DSD framework, this information or knowledge distills in a bottom-up manner on the encoder side and in a reverse top-down manner on the decoder side of a U-shaped backbone. Additionally, DSD also includes the distillation of knowledge from the ground-truth labels to the decoder layers of the U-shaped network, which is a process known as deep supervision in medical image segmentation [32]. The self-distillation process of DSD not only allows intermediate decoder and encoder representations to be optimized and aligned with the final representation of the U-shaped network, but also allows the hidden knowledge to flow from the deepest decoder and encoder layers to the shallower decoder and encoder layers respectively, and hence benefits from both regularization and generalization. In this manner, DSD leverages the benefits of deep supervision by overcoming optimization difficulties and achieving faster convergence [32].
@@
    We performed extensive experiments on various public 3D medical image segmentation datasets (one with cardiac substructures, one with brain tumors and another segmenting the Hippocampus), with DSD attached to several state-of-the-art U-shaped backbones (Vision and swin transformer based and CNN-based encoders) and demonstrated significant quantitative and qualitative improvements over those backbones with negligible increase in model parameters and training time.
A preliminary version of this work was presented at ISBI 2024 [33]. The rest of the paper is organized as follows: Section 2 describes the data set and the details of the dual self-distillation (DSD) framework. Section 3 describes the experimental setup and the results of our experiments. Section 4 presents discussion and future applications of our framework. Lastly, Section 5 presents our conclusion.
@@ 2.1 Datasets
We used two publicly available datasets to train and evaluate our proposed DSD framework.
Cardiac dataset (MMWHS) - This comprised of high resolution 3D CT angiography (CTA) data of 20 patients from the Multi-Modal Whole Heart Segmentation (MMWHS) dataset [46, 44, 45]. Each 3D image were of variable sizes along the three dimensions with a single input channel for the CT modality. The ground-truth labels consisted of 7 foreground classes of cardiac substructures and a background class. The foreground classes are: (i) Myocardium (MYO), (ii) Left Atrium (LA), (iii) Left Ventricle (LV), (iv) Right Atrium (RA), (v) Right Ventricle (RV), (vi) Ascending Aorta (AA), (vii) Pulmonary Artery (PA). We split the data into a training and validation set of 16 and 4 patients respectively. Then, we performed a 5-fold cross validation over the entire dataset.
We performed a set of preprocessing steps on the dataset prior to using them for training and evaluation with our network. The sequence of preprocessing steps used prior to training are: (i) We resampled all of the input images and ground-truth labels to have a voxel dimension of 1.5 mm along the coronal and sagittal direction and 2 mm along the axial direction. Bilinear interpolation was used to resample the images and nearest neighbor interpolation was used to resample the ground-truth labels. (ii) The orientation of all the input images and ground-truth labels were changed to have a uniform orientation of “Right”, “Anterior” and “Superior” for the coronal, sagittal and axial directions respectively. (iii) We scaled the intensity range of all the input images between 0 and 1 and then normalized each dimension of the image separately using the z-score normalization. (iv) To generate augmentations and have a uniform size for all the images and ground-truth labels, we randomly cropped out two 3D volumes of size (96,96,96) with the foreground voxel as the center. (v) These images and labels were then randomly flipped along all the three axis of the 3D volume with a probability of 10%. (vi) The images and labels were randomly rotated by 90 degrees with a probability of 10% along the coronal-sagittal plane. (vii) The intensity of the input images were randomly shifted between (-0.1,0.1) with a probability of 50%.
@@
Brain tumor dataset (MSD-BraTS) - The brain tumor segmentation (BraTS) task [40, 41, 39] was used from the Medical Segmentation Decathlon (MSD) dataset [43, 42]. This dataset comprised of a combination of BraTS 2016 and 2017 datasets. The segmentation task comprised of 484 patients having multi-modal multi-site MRI data. Each 3D image consisted of 4 channels for the multi-modal MRI input (FLAIR, T1w, T1gd and T2w). The ground-truth labels consisted of 3 foreground classes and a background class. The foreground classes are: (i) Whole tumor (WT), (ii) Enhancing tumor (ET) and (iii) Tumor core (TC). We took the approach adopted by UNETR [22] and split the data into a training, validation and testing set of 388, 72 and 24 patients respectively.
The sequence of preprocessing steps used prior to training are: (i) The orientation of all the input images and ground-truth labels were changed to have a uniform orientation of “Right”, “Anterior” and “Superior” for the coronal, sagittal and axial directions respectively. (ii) To generate augmentations and have a uniform size for all the images and ground-truth labels, we randomly cropped out two 3D volumes of size (128,128,128) with the foreground voxel as the center. (iii) These images and labels were then randomly flipped along all the three axis of the 3D volume with a probability of 50%. (iv) The images and labels were randomly rotated by 90 degrees with a probability of 10% along the coronal-sagittal plane. (v) The intensity of the input images were shifted between (-0.1,0.1) with a probability of 100%.
@@
Hippocampus Dataset (Hippocampus) - The Hippocampus segmentation task was used from the Medical Segmentation Decathlon (MSD) dataset [43, 42]. The segmentation task comprised of 260 patients with 3D T1-weighted MRI images. Each volume was annotated by using 2 labels for hippocampus and parts of the subiculum. Hence, the ground-truth labels consisted of 2 foreground classes and a background class. The foreground classes are: (i) Anterior, and (ii) Posterior. We divided this dataset into 200 patients for training and validation and 60 patients for held-out testing.
The sequence of preprocessing steps used prior to training are: (i) The orientation of all the input images and ground-truth labels were changed to have a uniform orientation of “Right”, “Anterior” and “Superior” for the coronal, sagittal and axial directions respectively. (ii) We scaled the intensity range of all the input images between 0 and 1. (iii) To generate augmentations and have a uniform size for all the images and ground-truth labels, we randomly cropped out 16 3D volumes of size (32,32,32) with the foreground voxel as the center. (iv) These images and labels were then randomly flipped along all the three axis of the 3D volume with a probability of 50%. (v) The images and labels were randomly rotated by 90 degrees with a probability of 10% along the coronal-sagittal plane. (vi) The intensity of the input images were shifted between (-0.1,0.1) with a probability of 100%.
@@
It should be noted that we are performing the random rotations only at 90-degree angles and not arbitrary angles. This is an important augmentation strategy since it simulates the CT or MRI acquisitions where the acquisition could be along any of the axial, coronal or sagittal planes. Hence, by randomly rotating the CT or MRI volumes by 90-degrees, we could inform the model about possible image acquisition strategies such that the model could generalize well during inference time to volumes that were acquired along the axial, coronal or sagittal planes.
@@
Figure 1: Self-distillation demonstrated with an U-shaped network for volumetric medical image segmentation. $Y$ and $G$ indicate the softmax of network output and one-hot encoded ground-truth (GT) labels, respectively. $E_{i}|_{i=1}^{Z}$ and $D_{i}|_{i=1}^{Z}$ denote the output of the bottleneck module on the encoder and decoder side, respectively. $T$ and $S$ denote the teacher and student probability distributions, respectively. All dashed lines shown are only used during training and removed during inference.
@@ 2.2 U-shaped backbone
The U-shaped backbone maps an input image $I$ ($I\in\mathbb{R}^{C\times H\times W\times D}$ with $H$, $W$, $D$ denoting the height, width and depth of a 3D input image, respectively and $C$ denoting number of imaging modalities/sequences such as CT or multi-modal MRI) to the ground-truth (GT) labels $G$ ($G\in\mathbb{R}^{K\times H\times W\times D}$ with $K$ classes).
@@
where, $G_{p,k}\in\mathbb{R}^{K\times H\times W\times D}$ denotes the one-hot encoded ground-truth labels and $Y_{p,k}\in\mathbb{R}^{K\times H\times W\times D}$ denotes the probability (softmax) output of the network for class $k$ at pixel $p$. $N=H*W*D$ denotes the total number of pixels in input image $I$ and ground-truth labels $G$. The Dice loss $L_{Dice}^{Y}$ measures the pixel-wise similarity while the cross-entropy loss $L_{CE}^{Y}$ measures the pixel-wise difference in distributions between the network output $Y$ and ground-truth labels $G$. This loss is back-propagated to the network to update the weights of the encoders and decoders.
@@ 2.4 U-shaped backbone with Dual Self-distillation (DSD)
We propose a novel dual self-distillation (DSD) framework for U-shaped backbones as shown by the purple and blue dashed arrows in Fig. 1. Our DSD framework consists of two main components.
(i) Distillation from ground-truth labels: the first part (purple dashed arrows in Fig. 1) is the distillation of knowledge from the ground-truth labels $G$ to each decoder of the U-shaped network. This process is known as deep supervision in medical image segmentation [32]. For our DSD framework, we calculate the Dice Cross-Entropy (DCE) loss (Eq. 1) between each decoder layer’s softmax output $D_{i}|_{i=1}^{Z}$ and the ground-truth labels $G$. This loss is defined as:
@@
where $D_{KL}(P^{S},P^{T})$ is the KL divergence between student ($P^{S}$) and teacher ($P^{T}$) probability distributions, $E_{i}$ and $D_{i}$ are the $i^{th}$ shallow encoder and decoder’s (student’s) softmax output ($P^{S}$) respectively, $E_{Z}$ and $D_{1}$ are the deepest encoder and decoder’s (teacher’s) softmax output ($P^{T}$), respectively, $Z$ is the number of encoders and decoders, $K$ denotes the number of classes of the ground-truth labels and $N$ denotes the total number of pixels.
We define our proposed dual self-distillation loss $L_{DSD}$ as follows:
@@
where $\alpha_{1}$ and $\alpha_{2}$ denote the coefficients that controls the amount of self-distillation between the encoder and decoder layers, respectively. Note that our generalized DSD framework is reduced to deep supervision when $\alpha_{1},\alpha_{2}=0$.
@@
where, $L_{DSD}$ is the dual self-distillation (DSD) loss and $\theta$ are the set of parameters that minimizes the DSD loss.

### v3

@@ Abstract
U-shaped networks and its variants have demonstrated exceptional results for medical image segmentation. In this paper, we propose a novel dual self-distillation (DSD) framework in U-shaped networks for volumetric medical image segmentation. DSD distills knowledge from the ground-truth segmentation labels to the decoder layers. Additionally, DSD also distills knowledge from the deepest decoder and encoder layer to the shallower decoder and encoder layers respectively of a single U-shaped network. DSD is a general training strategy that could be attached to the backbone architecture of any U-shaped network to further improve its segmentation performance. We attached DSD on several state-of-the-art U-shaped backbones, and extensive experiments on various public 3D medical image segmentation datasets (cardiac substructure, brain tumor and Hippocampus) demonstrated significant improvement over the same backbones without DSD. On average, after attaching DSD to the U-shaped backbones, we observed an increase of 2.82%, 4.53% and 1.3% in Dice similarity score, a decrease of 7.15 mm, 6.48 mm and 0.76 mm in the Hausdorff distance, for cardiac substructure, brain tumor and Hippocampus segmentation, respectively. These improvements were achieved with negligible increase in the number of trainable parameters and training time. Our proposed DSD framework also led to significant qualitative improvements for cardiac substructure, brain tumor and Hippocampus segmentation over the U-shaped backbones. The source code is publicly available at https://github.com/soumbane/DualSelfDistillation.
@@ 1 Introduction
The need of transferring knowledge from a large teacher network to a smaller student network was later eliminated by the process of self-distillation [30]. Self-distillation uses the deepest layer of a single model to act as the teacher network in order to distill knowledge to the shallower layers of the same model, acting as the student network. Since self-distillation uses a single model, it is much more efficient when compared to ensemble methods. Additionally, self-distillation is much faster to train as it is not necessary to train a large network as is required by knowledge distillation. Self-distillation has been applied for numerous computer vision tasks such as image classification [30, 35] and object detection [34]. A rigorous theoretical explanation of ensemble, knowledge distillation and self-distillation for image classification was recently provided in [47]. More recently, the self-distillation process was applied to a transformer based TransUNet [49], along with multi-scale fusion blocks to perform 3D medical image segmentation [36]. However, this method was limited to the specific TransUNet architecture.
@@
In this paper, we address the question of whether we can train a single U-shaped network and leverage the learned representations of the deeper encoder and decoder network layers to improve the prediction performance of medical image segmentation. Here, we propose a novel volumetric dual self-distillation (DSD) framework that could independently be attached to any U-shaped image segmentation backbone. In DSD, the deepest encoder and decoder of the U-shaped backbone act as the teacher networks for the shallower encoders and decoders which act as student networks. We found that the deepest encoder which is at the bottom of the contracting encoder path of the U-shaped network contains more contextual information compared to the shallower encoder layers. Similarly, the deepest decoder which is at the top of the expanding decoder path of the U-shaped network contains more semantic information than the shallower decoder layers. Thus, in our DSD framework, this information or knowledge distills in a bottom-up manner on the encoder side and in a reverse top-down manner on the decoder side of a U-shaped backbone. Additionally, DSD also includes the distillation of knowledge from the Ground-truth Labels (GT Labels) to the decoder layers of the U-shaped network, which is a process known as deep supervision in medical image segmentation [32]. The self-distillation process of DSD not only allows intermediate decoder and encoder representations to be optimized and aligned with the final representation of the U-shaped network, but also allows the hidden knowledge to flow from the deepest decoder and encoder layers to the shallower decoder and encoder layers respectively, and hence benefits from both regularization and generalization. In this manner, DSD leverages the benefits of deep supervision by overcoming optimization difficulties and achieving faster convergence [32].
@@
    We performed extensive experiments on various public 3D medical image segmentation datasets (one with cardiac substructures, one with brain tumors and another segmenting the Hippocampus), with DSD attached to several state-of-the-art U-shaped backbones (e.g., nnU-Net, UNETR, VNet and Swin UNETR) and demonstrated significant quantitative and qualitative improvements over those backbones with negligible increase in model parameters and training time.
A preliminary version of this work was presented at ISBI 2024 [33]. In this paper, we have thoroughly expanded the theoretical explanation of DSD as a regularization term, analyzed the effect of network parameters, and introduced new experiments with additional U-shaped networks. We also compared our framework with state-of-the-art methods and performed extensive ablation studies to validate our design choices. The rest of the paper is organized as follows: Section 2 describes the data set and the details of the DSD framework. Section 3 describes the experimental setup and the results of our experiments. Section 4 presents discussion and future applications of our framework. Lastly, Section 5 presents our conclusion.
@@ 2.1 Datasets
We used three publicly available datasets to train and evaluate our proposed DSD framework. These datasets were chosen as they cover different organs of the human body (whole heart, brain tumor and hippocampus), and they were acquired with different techniques such as Computed Tomography (CT), T1-weighted Magnetic Resonance Imaging (MRI) and multi-modal (T1w, T2w, FLAIR and T1c) multi-site MRI. Evaluating with these datasets would ensure that our proposed approach could be applied to various real-world radiology images involving different anatomies of the human body.
Figure 1: Self-distillation demonstrated with an U-shaped network for volumetric medical image segmentation. $Y$ and $G$ indicate the softmax of network output and one-hot encoded GT Labels, respectively. $E_{i}|_{i=1}^{Z}$ and $D_{i}|_{i=1}^{Z}$ denote the output of the bottleneck module on the encoder and decoder side, respectively. $T$ and $S$ denote the teacher and student probability distributions, respectively. All dashed lines shown are only used during training and removed during inference. The input is shown as a stack of images to indicate a 3D CT volume, and the output shows one-slice of the GT Labels overlayed on a 2D CT slice for a clear visualization of the different segmentation classes.
Cardiac dataset (MMWHS) - This comprised of high resolution 3D CT angiography (CTA) data of 20 patients from the Multi-Modal Whole Heart Segmentation (MMWHS) dataset [46, 44, 45]. Each 3D image were of variable sizes along the three dimensions with a single input channel for the CT modality. The GT Labels consisted of 7 foreground classes of cardiac substructures and a background class. The foreground classes are: (i) Myocardium (MYO), (ii) Left Atrium (LA), (iii) Left Ventricle (LV), (iv) Right Atrium (RA), (v) Right Ventricle (RV), (vi) Ascending Aorta (AA), (vii) Pulmonary Artery (PA). We split the data into a training and validation set of 16 and 4 patients respectively. Then, we performed a 5-fold cross validation over the entire dataset.
We performed a set of preprocessing steps on the dataset prior to using them for training and evaluation with our network. The sequence of preprocessing steps used prior to training are: (i) We resampled all of the input images and GT Labels to have a voxel dimension of 1.5 mm along the coronal and sagittal direction and 2 mm along the axial direction. Bilinear interpolation was used to resample the images and nearest neighbor interpolation was used to resample the GT Labels. (ii) The orientation of all the input images and GT Labels were changed to have a uniform orientation of “Right”, “Anterior” and “Superior” for the coronal, sagittal and axial directions respectively. (iii) We scaled the intensity range of all the input images between 0 and 1 and then normalized each dimension of the image separately using the z-score normalization. (iv) To generate augmentations and have a uniform size for all the images and GT Labels, we randomly cropped out two 3D volumes of size (96,96,96) with the foreground voxel as the center. (v) These images and labels were then randomly flipped along all the three axis of the 3D volume with a probability of 10%. (vi) The images and labels were randomly rotated by 90 degrees with a probability of 10% along the coronal-sagittal plane. (vii) The intensity of the input images were randomly shifted between (-0.1,0.1) with a probability of 50%.
@@
Brain tumor dataset (MSD-BraTS) - The brain tumor segmentation (BraTS) task [40, 41, 39] was used from the Medical Segmentation Decathlon (MSD) dataset [43, 42]. This dataset comprised of a combination of BraTS 2016 and 2017 datasets. The segmentation task comprised of 484 patients having multi-modal multi-site MRI data. Each 3D image consisted of 4 channels for the multi-modal MRI input (FLAIR, T1w, T1gd and T2w). The GT Labels consisted of 3 foreground classes and a background class. The foreground classes are: (i) Whole tumor (WT), (ii) Enhancing tumor (ET) and (iii) Tumor core (TC). We took the approach adopted by UNETR [22] and split the data into a training, validation and testing set of 388, 72 and 24 patients respectively.
The sequence of preprocessing steps used prior to training are: (i) The orientation of all the input images and GT Labels were changed to have a uniform orientation of “Right”, “Anterior” and “Superior” for the coronal, sagittal and axial directions respectively. (ii) To generate augmentations and have a uniform size for all the images and GT Labels, we randomly cropped out two 3D volumes of size (128,128,128) with the foreground voxel as the center. (iii) These images and labels were then randomly flipped along all the three axis of the 3D volume with a probability of 50%. (iv) The images and labels were randomly rotated by 90 degrees with a probability of 10% along the coronal-sagittal plane. (v) The intensity of the input images were shifted between (-0.1,0.1) with a probability of 100%.
@@
Hippocampus Dataset (Hippocampus) - The Hippocampus segmentation task was used from the Medical Segmentation Decathlon (MSD) dataset [43, 42]. The segmentation task comprised of 260 patients with 3D T1-weighted MRI images. Each volume was annotated by using 2 labels for hippocampus and parts of the subiculum. Hence, the GT Labels consisted of 2 foreground classes and a background class. The foreground classes are: (i) Anterior, and (ii) Posterior. We divided this dataset into 200 patients for training and validation and 60 patients for held-out testing.
The sequence of preprocessing steps used prior to training are: (i) The orientation of all the input images and GT Labels were changed to have a uniform orientation of “Right”, “Anterior” and “Superior” for the coronal, sagittal and axial directions respectively. (ii) We scaled the intensity range of all the input images between 0 and 1. (iii) To generate augmentations and have a uniform size for all the images and GT Labels, we randomly cropped out 16 3D volumes of size (32,32,32) with the foreground voxel as the center. (iv) These images and labels were then randomly flipped along all the three axis of the 3D volume with a probability of 50%. (v) The images and labels were randomly rotated by 90 degrees with a probability of 10% along the coronal-sagittal plane. (vi) The intensity of the input images were shifted between (-0.1,0.1) with a probability of 100%.
@@
It should be noted that during training, we are performing the random rotations only at 90-degree angles and not arbitrary angles. This is an important augmentation strategy since it simulates the CT or MRI acquisitions where the acquisition could be along any of the axial, coronal or sagittal planes. Hence, by randomly rotating the CT or MRI volumes by 90-degrees, we could inform the model about possible image acquisition strategies such that the model could generalize well during inference time to volumes that were acquired along the axial, coronal or sagittal planes. For evaluation purposes, the steps we performed are the standard steps followed by previous studies in the medical image segmentation literature ([20], [22], [23], [26]). These evaluation steps bring all volumes to the same orientation (“RAS” for our case) which is a crucial step necessary for all algorithms. Moreover, the uniform voxel dimension and scaled intensity are the least computationally intensive steps that could be easily performed on a new test image. Hence, all previous algorithms used them, and we followed the same approach in this paper for all baseline algorithms with and without DSD.
@@ 2.2 U-shaped backbone
The U-shaped backbone maps an input image $I$ ($I\in\mathbb{R}^{C\times H\times W\times D}$ with $H$, $W$, $D$ denoting the height, width and depth of a 3D input image, respectively and $C$ denoting number of imaging modalities/sequences such as CT or multi-modal MRI) to the GT Labels $G$ ($G\in\mathbb{R}^{K\times H\times W\times D}$ with $K$ classes).
@@
where, $G_{p,k}\in\mathbb{R}^{K\times H\times W\times D}$ denotes the one-hot encoded GT Labels and $Y_{p,k}\in\mathbb{R}^{K\times H\times W\times D}$ denotes the probability (softmax) output of the network for class $k$ at pixel $p$ (here $p$ denotes the 3D pixel position, i.e., one 3D coordinate of ($H,W,D$)). $N=H*W*D$ denotes the total number of pixels in input image $I$ and GT Labels $G$. The Dice loss $L_{Dice}^{Y}$ measures the pixel-wise similarity while the cross-entropy loss $L_{CE}^{Y}$ measures the pixel-wise difference in distributions between the network output $Y$ and GT Labels $G$. This loss is back-propagated to the network to update the weights of the encoders and decoders.
@@ 2.4 U-shaped backbone with Dual Self-distillation (DSD)
We propose a novel DSD framework for U-shaped backbones as shown by the purple and blue dashed arrows in Fig. 1. Our DSD framework consists of two main components.
(i) Distillation from GT Labels: the first part (purple dashed arrows in Fig. 1) is the distillation of knowledge from the GT Labels $G$ to each decoder of the U-shaped network. This process is known as deep supervision in medical image segmentation [32]. For our DSD framework, we calculate the Dice Cross-Entropy (DCE) loss (Eq. 1) between each decoder layer’s softmax output $D_{i}|_{i=1}^{Z}$ and the GT Labels $G$. This loss is defined as:
@@
where $D_{KL}(P^{S},P^{T})$ is the KL divergence between student ($P^{S}$) and teacher ($P^{T}$) probability distributions, $E_{i}$ and $D_{i}$ are the $i^{th}$ shallow encoder and decoder’s (student’s) softmax output ($P^{S}$) respectively, $E_{Z}$ and $D_{1}$ are the deepest encoder and decoder’s (teacher’s) softmax output ($P^{T}$), respectively, $Z$ is the number of encoders and decoders, $K$ denotes the number of classes of the GT Labels and $N$ denotes the total number of pixels.
We define our proposed DSD loss $L_{DSD}$ as follows:
@@
where $\alpha_{1}$ and $\alpha_{2}$ denote the coefficients that controls the amount of self-distillation between the encoder and decoder layers, respectively. Note that our DSD framework is reduced to deep supervision when $\alpha_{1},\alpha_{2}=0$.
@@
where, $L_{DSD}$ is the DSD loss and $\theta$ are the set of parameters that minimizes the DSD loss.